# What Makes a Visualization Complex?

### v1

@@ Appendix D Additional Implications on Using Complexity to Support Design
Ideally, this project outcome would support new design. Declutter removes any additional ink not needed: grid lines, labels, colors, and 3D effects are taken as redundant visual elements [2]. The color removal in their decluttering experiment was appropriate for conditions where colors were not used to encode data. Subsequently, the authors added highlight colors and text annotations to convey key takeaways from the figures. In doing so, the purpose of the images shifted from classic rhetorical visualizations to communicative tools, integrating textual explanations with semiotic marks. The authors observed a weak improvement in perceived VC, a finding that aligns with insights from our inquiry. On the other hand, human perception of “redundancy” depends on viewers’ expertise: people with less expertise would be benefited by more data-relevant elements.
Our regression model revealed that metrics associated with edge density and feature points, along with the color count, play a dominant role in determining perceived complexity, which aligns closely with the work of Oliva et al. [99]. For visualization applications, extensive studies on visual abstraction and suggestive contours, which are driven by perceptual principles [138, 30], could potentially help reduce VC.
| O.IE | ||||||||||||
| O.KC | 0.001 | |||||||||||
| O.SE | 0.001 | 0.001 | ||||||||||
| O.IG | 0.001 | 0.001 | 0.001 | |||||||||
| O.FC | 0.001 | 0.001 | 0.001 | 0.001 | ||||||||
| O.H | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | |||||||
| O.CF | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | ||||||
| O.ERGB | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | |||||
| O.ED | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | 0.001 | ||||
@@
| Entropy | The amount of information required to encode the distribution of pixel values. | [124, 122] | To measure the information content and "visual richness" of an image. | |
@@
| Noise | The unwanted variations and grains present in the image. | [77] | To propose a 3D feature space to represent visual complexity of images. | |
@@
| Material complexity | The richness, variety and intricacy of materials and textures. | [110] | To measure the perceived visual complexity in a computer graphics scene. | |
@@
| Number of objects | The number of objects or components that make up a visual scene. | [99] | To evaluates clutter measurement methods of perceived scene complexity. | |

### v3

@@ Appendix D Additional Implications on Using Complexity to Support Design
Ideally, this project outcome would support new design. Declutter removes any additional ink not needed: grid lines, labels, colors, and 3D effects are taken as redundant visual elements [2]. The color removal in their decluttering experiment was appropriate for conditions where colors were not used to encode data. Subsequently, the authors added highlight colors and text annotations to convey key takeaways from the figures. In doing so, the purpose of the images shifted from classic rhetorical visualizations to communicative tools, integrating textual explanations with semiotic marks. The authors observed a weak improvement in perceived VC, a finding that is consistent with the insights from our inquiry. On the other hand, human perception of “redundancy” depends on the expertise of the viewer: People with less experience would benefit from more data-relevant elements.
Our regression model revealed that metrics associated with edge density and feature points, along with color count, play a dominant role in determining perceived complexity, which is in close alignment with the work of Oliva et al. [98]. For visualization applications, extensive studies of visual abstraction and suggestive contours, which are driven by perceptual principles [140, 30], could potentially help reduce VC.
@@
| Entropy | The amount of information required to encode the distribution of pixel values. | [124, 122] | To measure the information content and “visual richness” of an image. | |
@@
| Noise | The unwanted variations and features present in the image. | [75] | To propose a 3D feature space to represent visual complexity of images. | |
@@
| Material complexity | The richness, variety, and intricacy of materials and textures. | [110] | To measure the perceived visual complexity in a computer graphics scene. | |
@@
| Number of objects | The number of objects or components that make up a visual scene. | [98] | To evaluate clutter. | |